Describe the translations applied to the graph of y 5 x2 to obtain a K graph of each quadratic function. Sketch the graph.

y=(x+1)^2-2
1
Expert's answer
2020-12-03T13:49:23-0500

First, we can compress the initial function "y=5x^2" along y-axis. We get


"y=x^2."


Then we can shift the function by 2 units down y-axis:


"y=x^2-2."


Then we shift the function left 1 unit and here is what we get:


"y=(x+1)^2-2."

In graph below, red is the initial function, purple is the final function: